Find the minimum sample size required to estimate population proportion. Margin Error: .09;confidence level: 90%; from a prior study, population proportion is estimated by .23
Answer by jim_thompson5910(35256)

use the formula n = p(1-p)(z/E)^2 where


n = min sample size
p = proportion (use p = 0.5 if this isn't given)
z = critical value (based on the confidence level)
E = margin of error


In this case, we know


p = 0.23
z = 1.645 (found using a table like this)
E = 0.09


Plug these values into the formula to get


n = p(1-p)(z/E)^2

n = 0.23*(1-0.23)*(1.645/0.09)^2

n = 59.1650651234568

n = 60 ... Always round UP to the nearest whole number.


So your sample size needs to be at least 60.